OpenStudy (anonymous):

how to find the derivative of 3^(2^(x)^2-3x)

myininaya (myininaya):

\[3^{2^{x^2}-3x}\] this?

OpenStudy (anonymous):

yes

myininaya (myininaya):

\[y=\frac{3^{2^{x^2}}}{3^{3x}} \text{ First I'm going to rewrite this using law of exponents }\] Now I will take ln( ) of both sides : \[\ln(y)=\ln(3^{2^{x^2}})-\ln(3^{3x})\] Now I'm going to use power rule for log... \[ln(y)=2^{x^2} \ln(3)-3x \ln(3)\] Now differentiate both sodes... We may find that we have to find the derivative of 2^(x^2) on some scratch paper by itself.

OpenStudy (anonymous):

ok how do you do that?

myininaya (myininaya):

using ln( ) on both sides then differentiating both sides

OpenStudy (anonymous):

how do you differentiate on both sides?

myininaya (myininaya):

\[u=2^{x^2} \\ \ln(u)=x^2 \ln(2) \text{ now differentiate }\]

myininaya (myininaya):

well you should know the derivative of x^2 is 2x and the derivative of ln(u) is u'/u

OpenStudy (anonymous):

well yeah but why didn't u take the derivative of 2x^2

OpenStudy (anonymous):

its not 2 to the power of x squared. It's 2x^2

myininaya (myininaya):

So are you telling me the problem I posted above was not the correct interpretation of the problem you posted?

OpenStudy (anonymous):

yeah I'm sorry. That was my fault. i didn't realize you put 2 to the power of x^2. it's just supposed to be 2x^2

myininaya (myininaya):

\[3^{2x^2-3x} ?\]

OpenStudy (anonymous):

yes that is correct

myininaya (myininaya):

\[y=3^{2x^2-3x} \\ \text{ still you need to take } \ln( ) \text{ of both sides } \\ \ln(y)=(2x^2-3x)\ln(3)\] now just differentiate both sides (this problem is a lot easier than the one before since we have less powers)

myininaya (myininaya):

\[(\ln(y))'=(2x^2-3x)'\ln(3)\]

OpenStudy (anonymous):

okay so next i take the derivative of (2x^2-3x) ?

myininaya (myininaya):

well you take the derivative of both sides

myininaya (myininaya):

the derivative of ln(y) =? the derivative of (2x^2-3x)=?

OpenStudy (anonymous):

uhm so 4x-3..

myininaya (myininaya):

\[(\ln(y))'=(2x^2-3x)'\ln(3) \\ (\ln(y))'=(4x-3)\ln(3)\] still need to differentiate the left hand side

OpenStudy (anonymous):

uhm im not quite sure how to do that..

myininaya (myininaya):

derivative (w.r.t. x) of ln(x) is 1/x derivative of ln(u) (w.r.t x) is u'/u so here you have y's instead of u's

myininaya (myininaya):

are you still here

myininaya (myininaya):

derivative of ln(y) (w.r.t x ) is just y'/y

myininaya (myininaya):

\[\frac{y'}{y}=(4x-3) \ln(3)\]

myininaya (myininaya):

solve for y'

myininaya (myininaya):

and then replace y with what you started with (since that is what we called y)

OpenStudy (anonymous):

okay so the answer is y'/y= (4x-3)ln(3) ?

myininaya (myininaya):

well have you solve for y' yet? and replace y with what we called y in the beginning because that is what the lines above your post say to do...

myininaya (myininaya):

look here is an example: \[u=5^x \\ \ln(u)=\ln(5^x) \\ \ln(u)=x \ln(5) \\ \frac{u'}{u}=1 \ln(5) \\ \frac{u'}{u}=\ln(5) \\ u'=u \ln(5) \\ u' = 5^x \ln(5) \]

myininaya (myininaya):

to solve for u' I multiplied both sides by u

myininaya (myininaya):

u was called 5^x so I replaced u with 5^x

myininaya (myininaya):

here is another example \[v=5^{x^2-3x+1} \\ \ln(v)=(x^2-3x+1)\ln(5) \\ \frac{v'}{v}=(2x-3)\ln(5) \\ v'=(2x-3) \ln(5) v \\ v'=(2x-3)\ln(5) 5^{x^2-3x+1}\]

OpenStudy (anonymous):

so y'= 3^(2x^2-3x)(4x-3)(ln3) ?

myininaya (myininaya):

sounds great

OpenStudy (anonymous):

thank you so much! So i have another problem that I'm stuck on as well. Do you mind guiding me through that one? I understand if you're too busy but if you can i'll really appreciate it.

myininaya (myininaya):

i have to eat sorry

OpenStudy (anonymous):

oh okay. well thanks anyways:)